# How Temporal Unrolling Supports Neural Physics Simulators

### v1

@@ Abstract
Unrolling training trajectories over time strongly influences the inference accuracy of neural network-augmented physics simulators. We analyze these effects by studying three variants of training neural networks on discrete ground truth trajectories. In addition to commonly used one-step setups and fully differentiable unrolling, we include a third, less widely used variant: unrolling without temporal gradients. Comparing networks trained with these three modalities makes it possible to disentangle the two dominant effects of unrolling, training distribution shift and long-term gradients. We present a detailed study across physical systems, network sizes, network architectures, training setups, and test scenarios. It provides an empirical basis for our main findings: A non-differentiable but unrolled training setup supported by a numerical solver can yield 4.5-fold improvements over a fully differentiable prediction setup that does not utilize this solver. We also quantify a difference in the accuracy of models trained in a fully differentiable setup compared to their non-differentiable counterparts. While differentiable setups perform best, the accuracy of unrolling without temporal gradients comes comparatively close. Furthermore, we empirically show that these behaviors are invariant to changes in the underlying physical system, the network architecture and size, and the numerical scheme. These results motivate integrating non-differentiable numerical simulators into training setups even if full differentiability is unavailable. We also observe that the convergence rate of common neural architectures is low compared to numerical algorithms. This encourages the use of hybrid approaches combining neural and numerical algorithms to utilize the benefits of both.
@@ 1 Introduction
Our understanding of physical systems relies on capturing their dynamics in mathematical models, often representing them with a partial differential equation (PDE). Forecasting the behavior with these models thus involves the notoriously costly and difficult task of solving the PDE. By aiming to increase simulator efficiencies, machine learning was successfully deployed to augment traditional numerical methods for solving these equations. Common areas of research are network architectures [51, 33, 19, 57], reduced order representations [38, 65, 16, 10, 68], and training methods [56, 52, 39, 7]. Previous studies were further motivated by performance and accuracy boosts offered by these methods, especially on GPU architectures. Especially computational fluid dynamics represents an application of machine learning augmented solvers [4, 13]. Potential applications in turbulence include closure modeling in Reynolds-averaged Navier-Stokes modeling [35, 71, 72] as well as closure models for large eddy simulation [43, 2, 37, 52, 58].
@@
Out of all physical systems of interest in simulations, those exhibiting transient dynamics are particularly challenging to model. Solutions to these systems are typically based on a time-discrete evolution of the underlying dynamics. Consequently, the simulator of such dynamics is subject to many auto-regressive invocations. Designing simulators that achieve stability and accuracy over long auto-regressive horizons can be challenging in numerical and neural settings. However, unlike numerical approaches, where unstable components such as dispersion errors are well studied and effective stabilizing approaches exist, little is known about the origin of instabilities and mitigating strategies for neural networks. As an interesting development, a stability analysis of echo-state networks was presented recently [40], allowing deeper insights into the behavior of recurrent networks. Nevertheless, stabilizing autoregressive neural simulators remains particularly difficult. One promising strategy is unrolling discrete trajectories at training time, improving long-term stability in different settings [56, 28, 52]. It is now a common approach when training neural networks that predict PDE trajectories on their own, or support numerical solvers in doing so. However, we still lack a systemic understanding of unrolling and its effects on network training.
@@
Within this paper, we explore unrolling from a dynamical systems perspective. A first fundamental insight is that no machine-learned approximation identically fits the underlying target. For physics simulators, this means that the dynamics of the learned system will always differ from the target dynamics. As such, the attractors for the target and the learned system do not coincide entirely. Unrolling can be seen as an exploration of the machine-learned dynamics’ attractor at training time. In contrast to training without unrolling, this approach thus allows sampling from a different distribution in the solution space. An unrolled training procedure samples from the learned dynamics and exposes (more of) the inference attractor at training time. As such, this difference between training and inference, commonly known as data shift in the machine learning community, is reduced. For longer horizons, more of the learned dynamics’ attractor is exposed, stabilizing the neural simulator.
@@
A central and rarely used variant employs unrolling without backpropagating gradients through the simulator. We study this training variant in detail because of its high practical relevance, as it can be realized by integrating existing, non-differentiable solvers into a deep learning framework. The forward pass of a non-differentiable unrolling setup is identical to the fully differentiable setup and thus features the same attractor-fitting property. In the backward pass, gradients are set to zero whenever they need to be propagated through a numerical solver. This effectively truncates the backpropagation, and avoids long gradient propagation chains. While cutting gradients potentially has a stabilizing effect, we find that it also limits the effective unrolling horizon, albeit for different reasons: the computed gradient approximations diverge from the true gradient as the unrolling horizon increases, making the learning process less reliable. Figure 1 illustrates these gradient inaccuracies for both training modalities. To the best of our knowledge, extensive studies of non-differentiable unrolled training do not yet exist. This is despite its high relevance to the scientific computing community as an approachable variant integrating numerical solvers in network training.
@@
We analyze these two unrolling variants using standardized solver and network architectures, while contrasting them with one-step training. This allows disentangling the effects of training data shift introduced by forward unrolling and accurate long-term gradients calculated by backpropagation. We test our findings on various physical systems with multiple network architectures, including convolutional and graph networks. Five recommendations for unrolled neural simulators training are derived from our results:
(I) Non-differentiable unrolling as an alternative
The mitigating effects on data shift introduced by unrolling can boost model performance. An unrolled but non-differentiable chain can increase performance by 23% on average. In addition, non-differentiable neural hybrid simulators do not require a new implementation of existing numerical solvers and thus represent an attractive alternative for fields with large traditional code bases. When no differentiable implementations are available, previous work typically resorts to pure prediction models. Instead, our results strongly encourage interfacing non-differentiable numerical implementations with machine learning, as doing so yields up to 4.5-fold accuracy improvements over prediction setups for identical network architectures and sizes.
(II) Differentiable unrolling delivers the best accuracy
When other hyperparameters are fixed, a differentiable unrolling strategy consistently outperforms other setups. At the cost of increased software engineering efforts, differentiable unrolling outperforms its one-step counterpart by 38% on average. Long-term gradients prove to be especially important for hybrid setups where neural networks correct numerical solvers.
(III) Hybrid approaches yield improved scaling:
Our study across network sizes finds them predominant regarding accuracy on inter- and extrapolative tests. However, we observe suboptimal convergence of the inference error when increasing the parameter count. These convergence rates compare poorly to the ones usually observed for numerical solvers. Most neural PDE simulators directly compete with numerical solvers. Thus, scaling to larger problems is critical for neural simulators. Unrolling can effectively train hybrid simulators even when integrated with a non-differentiable solver. These combine the benefits of neural networks with the scaling properties of numerical solvers.
(IV) Low-dimensional problems match
(V) Curriculums are necessary
The training of unrolled setups is non-trivial and sensitive to hyperparameters. Reliable training setups utilize a curriculum where the unrolled number of steps slowly increases. The learning rate needs to be adjusted to keep the amplitude of the gradient feedback at a stable level. Our work takes a task- and architecture-agnostic stance. Combined with the large scale of evaluations across more than three thousand models, this allows for extracting general trends and the aforementioned set of recommendations. Source code and training data will be made available upon acceptance.
2 Related work
Data-driven PDE solvers:
Machine learning based simulators aim to address the limitations of classical PDE solvers in challenging scenarios with complex dynamics [18, 60, 9]. While physics-informed networks have gained popularity in continous PDE modeling [48, 13], many learned approaches work in a purely data-driven manner on discrete trajectories. Several of these use advanced network architectures to model the time-evolution of the PDE, such as graph networks [46, 51], problem-tailored architectures [61, 53], bayesian networks [70], generative adverserial networks [67], transformer models [22, 20, 32], or lately diffusion models [34, 36, 29]. Variations of these approaches compute the evolution in an encoded latent space [65, 19, 20, 10, 68, 59].
Unrolled training:
Fitting an unrolled ground truth trajectory is frequently used for training autoregressive methods. This can be done with unrolled architectures that solely rely on networks [19], or for networks that correct solvers [56, 28, 39, 41]. For the latter, differentiable or adjoint solvers allow backpropagation of gradients through the entire chain [25, 52, 24]. The studies building on these solvers report improved network performance for longer optimization horizons. However, differentiability is rarely satisfied in existing code bases. A resulting open question is how much the numerical solver’s differentiability assists in training accurate networks. As the introduction mentions, two properties of differentiable unrolling affect the training procedure. The data shift moves the observed training data closer toward realistic inference scenarios [66, 62]. For instance, [7, 47] proposed variations of classic truncation [55], and reported a positive effect on the network performance, where ”warm-up” steps without contribution to the learning signal stabilize the trained networks. The differentiability of the numerical solver introduces the second property, which allows the propagation of gradients through time evolutions. The resulting loss landscape better approximates temporal extrapolation, which benefits model performance [56, 52, 28]. Recently, possible downsides of differentiable unrolling concerning gradient stability were investigated. While a vanishing/exploding effect is well known for recurrent networks, it is only sparsely studied for hybrid approaches. It was found that the stability of the backpropagation gradients aligns with the Lyapunov time of the physical system for recurrent prediction networks [42]. As a potential remedy, it was proposed to cut the backpropagation chain into individual sequences [37].
Benchmarks and datasets:
Several benchmarks and datasets have been published to increase comparability and promote standardization of machine-learned PDE simulators, especially in fluid mechanics. A dataset of measured real-world smoke clouds can be found in [14]. A high-fidelty Reynolds-averaged Navier-Stokes dataset is provided in [6]. Furthermore, more specialized datasets like [69] focus on fluid manipulation and robotics. A dataset of large-scale $2$D fluid flows on non-uniform meshes and a benchmark for transformer models are also available [27]. As we focus on tasks that correct solvers, we evaluate our models on frameworks with differentiable solvers [24, 56].
@@ 3 Unrolling PDE Evolutions
Unrolling is a common strategy for learning time sequences. A first important distinction is whether the task at hand is a prediction or a correction task. Additionally, we formally introduce the differences between non-differentiable and differentiable unrolling in gradient calculation. These are then used to derive an intuition about the gradient effects of unrolling.
@@ 3.1 Evolving partial differential equations with neural networks
with $\mathbf{u}$ representing the field variables of the physical system. The dynamics of $\mathcal{F}$ have an attractor $A_{\mathcal{F}}$ representing a subspace of the solution space. A state contained in $A_{\mathcal{F}}$ will always evolve to other states on $A_{\mathcal{F}}$, while states in the neighborhood of $A_{\mathcal{F}}$ converge to this attractor. We study Neural Simulator architectures that use neural networks for evolving discretized forms of PDE s. They are autoregressively applied to generate trajectories of discrete $\mathbf{u}^{t}$. Prediction setups fully rely on a neural network to calculate the next timestep as $\mathbf{u}^{t+1}=f_{\theta}(\mathbf{u}^{t})$, where $\theta$ represents the parameters of the network function. In a prediction configuration, the network fully replaces a numerical solver with the goal of improving its accuracy and performance. Similarly, correction setups are also concerned with time-evolving a discretized PDE but additionally include a physical prior $\mathcal{S}$ that approximates the solution. The network then corrects this approximation such that $\mathbf{u}^{t+1}=f_{\theta}(\mathcal{S}(\mathbf{u}^{t}))$. These architectures represent a hybrid between numerical and neural architectures, potentially combining the advantages of both domains. Possible applications include, turbulence modeling in RANS or LES, as well as dynamical system control setups.
Crucially, inputs and outputs of the networks live in the space of the solution vector field, both for predictive and corrective setups. Thus, we use different instances of the same architecture in both setups. We use the same notation $f_{\theta}$ for both applications to stress this property.

### v2

# Differentiability in Unrolled Training of Neural Physics Simulators on Transient Dynamics

@@ Abstract
Unrolling training trajectories over time strongly influences the inference accuracy of neural network-augmented physics simulators. We analyze this in three variants of training neural time-steppers. In addition to one-step setups and fully differentiable unrolling, we include a third, less widely used variant: unrolling without temporal gradients. Comparing networks trained with these three modalities disentangles the two dominant effects of unrolling, training distribution shift and long-term gradients. We present detailed study across physical systems, network sizes and architectures, training setups, and test scenarios. It also encompasses two simulation modes: In prediction setups, we rely solely on neural networks to compute a trajectory. In contrast, correction setups include a numerical solver that is supported by a neural network. Spanning these variations, our study provides the empirical basis for our main findings: Non-differentiable but unrolled training with a numerical solver in a correction setup can yield substantial improvements over a fully differentiable prediction setup not utilizing this solver. The accuracy of models trained in a fully differentiable setup differs compared to their non-differentiable counterparts. Differentiable ones perform best in a comparison among correction networks as well as among prediction setups. For both, the accuracy of non-differentiable unrolling comes close. Furthermore, we show that these behaviors are invariant to the physical system, the network architecture and size, and the numerical scheme. These results motivate integrating non-differentiable numerical simulators into training setups even if full differentiability is unavailable. We show the convergence rate of common architectures to be low compared to numerical algorithms. This motivates correction setups combining neural and numerical parts which utilize benefits of both.
@@ 1 Introduction
Our understanding of physical systems relies on capturing their dynamics in mathematical models, often representing them with a partial differential equation (PDE). Forecasting the behavior with these models thus involves the notoriously costly and difficult task of solving the PDE. By aiming to increase simulator efficiencies, machine learning was successfully deployed to augment traditional numerical methods for solving these equations. Common areas of research are network architectures [79, 44, 27, 94, 91], reduced order representations [51, 104, 3, 59, 14, 107], and training methods [93, 83, 52, 10]. Previous studies were further motivated by performance and accuracy boosts offered by these methods, especially on GPU architectures. Machine learning based simulators aim to address the limitations of classical PDE solvers in challenging scenarios with complex dynamics [26, 98, 13]. Computational fluid dynamics represents an application of machine learning augmented solvers [7, 21]. This especially includes turbulence closure modeling in Reynolds-averaged Navier-Stokes modeling [48, 114, 115, 112, 113] as well as closure models for large eddy simulation [60, 5, 50, 83, 95]. In tackling these challenges, various architectural advancements have recently been made. On the one hand, physics-informed networks have gained popularity in continuous PDE modeling [73, 21, 87], and extensions of this have been made to discrete representations [74, 16]. However, many learned approaches work in a purely data-driven manner on discrete trajectories. Several of these use advanced network architectures to model the time-evolution of the PDE, such as graph networks [66, 79] and particle simulators [103, 70], problem-tailored architectures [99, 84, 19], bayesian networks [109], generative adverserial networks [106], transformer models [32, 28, 42], or lately diffusion models [47, 49, 38]. Variations of these approaches compute the evolution in an encoded latent space [104, 27, 28, 14, 107, 96]. In an effort to achieve discretization-independence of the learned simulators, neural operators have shifted the learning task towards finding a mapping between the function spaces of the numerical solutions. The architectural changes that become necessary when moving to this paradigm have been made for convolutional architectures [75], graph networks [45] or Fourier-based features [44], among others. Additionally, physical knowledge is often integrated into the neural solver by keeping some classical numerical component as part of the solver architecture [61, 110]. In this context, the practice of training neural networks in combination with numerical solvers emerged. The resulting simulator is sometimes referred to as hybrid solver or augmented solver in literature [37, 50, 49]. Since neural network training is based on automatic differentiation, differentiable solvers are required to calculate network updates. In his setting, network optimization is conceptually equivalent to adjoint optimization. Here, the adjoint equation provides ”sensitivities”, which are referred to as ”gradients” in the discrete automatic differentiation setting [80]. Due to this similarity, early works utilized adjoint solvers into which networks were embedded [83, 52].
@@
Out of all physical systems of interest in simulations, those exhibiting transient dynamics are particularly challenging to model. Solutions to these systems are typically based on a time-discrete evolution of the underlying dynamics. Consequently, the simulator of such dynamics is subject to many auto-regressive invocations. Designing simulators that achieve stability and accuracy over long auto-regressive horizons can be challenging in numerical and neural settings. However, unlike numerical approaches, where unstable components such as dispersion errors are well studied and effective stabilizing approaches exist, little is known about the origin of instabilities and mitigating strategies for neural networks. As an interesting development, a stability analysis of echo-state networks was presented recently [54], allowing deeper insights into the behavior of recurrent networks. Nevertheless, stabilizing autoregressive neural simulators remains particularly difficult.
One promising strategy is unrolling, where multi-step trajectories are evolved at training time. Um et al. [93] used an auto-differentiable implementation of an incompressible Navier-Stokes solver to train correction networks to accelerate numerical simulations of a cylinder-wake flow. Kochkov et al. [37] used a very similar technique also employing unrolled correction learning on the two-dimensional Kolmogorov flow. Both of these setups used neural network architectures based on discrete convolutions. An adjoint solver was used by Sirignano et al. [83] on challenging three-dimensional turbulence scenarios to train a point-wise closure model for Large-Eddy simulation. These previous works all successfully deployed unrolling. In this practice, one update to the network weights is based on an unrolled trajectory, where the neural simulator is applied multiple times in an auto-regressive fashion. By design, these unrolled training trajectories are closer to the inference scenario than simple one-step training. This effectively reduces the data shift, which describes the difference in the distribution of training and inference data. By exposing autoregressive trajectories at training time, unrolling moves the observed training data closer toward realistic inference scenarios [105, 100]. A measurable gain in long-term stability of networks trained with this method was found in previous work [93, 50, 55]. By now, unrolling represents a widely used approach when training neural networks that predict PDE trajectories on their own [27], or support numerical solvers in doing so, for instance, in LES closure modeling [83, 52, 50]. An important observation is that these previous works utilized the full backpropagation gradient for network optimization. In an unrolled setup including a numerical solver, this can only be calculated using differentiable or adjoint solvers. Since differentiability is rarely satisfied in existing code bases, this is a first limitation of differentiable unrolling. A resulting open question is how much the numerical solver’s differentiability assists in training accurate networks.
Additionally, possible downsides of differentiable unrolling concerning gradient stability were investigated in recent work. While a vanishing/exploding effect is well known for recurrent networks [31], it is only sparsely studied for correction approaches featuring numerical and neural components. The stability of the backpropagation gradients was found to align with the Lyapunov time of the physical system for recurrent prediction networks [56]. Furthermore, unrolling for too many steps was associated with diminishing returns [93, 50] and even showed negative effects in previous work [55]. This indicates that the gradient stability of unrolled setups dependent on the unrolled horizon marks an important limiting factor. In practical considerations, unrolled setups need to balance the benefits of data shift reduction with sufficient gradient stability. The reported success of some methods used in previous work can be related to this balancing act. For instance, recent studies [10, 70] proposed a variant of classic truncated backpropagation in time [89]. These recent works reported a positive effect on the network performance when ”warm-up” steps were introduced. A warm-up step increases the unrolled horizon without contributing to the learning signal since backpropagation gradients are discarded for these steps. This effect was reported to stabilize the trained networks. As a similar potential remedy, it was proposed to cut the backpropagation chain into individual sequences [50].
@@
Within this paper, we explore unrolling from a dynamical systems perspective. A first fundamental insight is that no machine-learned approximation identically fits the underlying target. For physics simulators, this means that the dynamics of the learned system will always differ from the target dynamics. As such, the attractors for the target and the learned system do not coincide entirely. Unrolling can be seen as an exploration of the machine-learned dynamics’ attractor at training time. In contrast to training without unrolling, this approach thus allows sampling from a different distribution in the solution space. An unrolled training procedure samples from the learned dynamics and exposes (more of) the inference attractor at training time. The difference between training and inference, which we referred to as data shift, is reduced. For longer horizons, more of the learned dynamics’ attractor is exposed, stabilizing the neural simulator.
@@
A central and rarely used variant employs unrolling without backpropagating gradients through the simulator. We study this training variant in detail because of its high practical relevance, as it can be realized by integrating existing, non-differentiable solvers into a deep learning framework. The forward pass of a non-differentiable unrolling setup is identical to the fully differentiable setup and thus features the same attractor-fitting property. In the backward pass, the non-differentiability of the solver is equivalent to an operator returning a zero gradient. This effectively truncates the backpropagation and avoids long gradient propagation chains. While cutting gradients potentially has a stabilizing effect, we find that it also limits the effective unrolling horizon, albeit for different reasons: the computed gradient approximations diverge from the true gradient as the unrolling horizon increases, making the learning process less reliable. Figure 1 illustrates these gradient inaccuracies for both training modalities. To the best of our knowledge, extensive studies of non-differentiable unrolled training do not yet exist. This is despite its high relevance to the scientific computing community as an approachable variant integrating numerical solvers in network training.
@@
We analyze these two unrolling variants using standardized solver and network architectures while contrasting them with one-step training. This allows disentangling the effects of data shift introduced by forward unrolling and accurate long-term gradients calculated by backpropagation. Our study targets four open questions in the context of unrolling:
(I) How effective is non-differentiable unrolling?
The vast majority of existing numerical solvers are not differentiable. We investigate the effectiveness of non-differentiable unrolling and whether integrating such a solver into the training loop yields measurable benefits. While previous studies have provided a good intuition on the effectiveness of differentiable unrolling, there are no prior results on non-differentiable unrolling.
(II) How do the unrolling variants compare?
Differentiable unrolling, non-differentiable unrolling, and one-step learning require different levels of software engineering efforts. A comparison of these training setups is necessary to choose the most suitable setup for a given application.
(III) Which benefits does correction training yield?
Whether machine learning architectures should be trained to augment numerical codes in a corrective setup or rather constitute surrogates in a predictive setup, is up to scientific debate. The integration of numerical solvers as indicated in (I) and (II) has to be motivated by distinct benefits of such correction approaches.
(IV) Do the effects of unrolling show in different PDE s?
The concepts of unrolling and differentiablility may arise in various applications governed by different physical laws. As such, we want to investigate how the effects of unrolling translate between physical systems, solvers and network architectures.
We test our findings on various physical systems with multiple network architectures, including convolutional and graph networks. To achieve this, we develop differentiable solvers that could serve as a benchmark in future studies. In this context, it is worth noting that several benchmarks and datasets have been published to increase comparability and promote standardization of machine-learned PDE simulators, especially in fluid mechanics. Datasets for advection-diffusion type PDEs [90], and measured real-world smoke clouds [22] have been proposed. A high-fidelity Reynolds-averaged Navier-Stokes dataset is provided in [9]. Furthermore, more specialized datasets focus on fluid manipulation [108], and navigation for aerial vehicles [36]. We choose three widely used differentiable scenarios and an additional non-differentiable case as the basis for the following investigation.
@@ 2 Unrolling PDE Evolutions
Unrolling is a common strategy for learning time sequences [93, 52, 37, 50]. A first important distinction is whether the task at hand is a prediction or a correction task. Additionally, we formally introduce the differences between non-differentiable and differentiable unrolling in gradient calculation. These are then used to derive an intuition about the gradient effects of unrolling.
@@ 2.1 Evolving partial differential equations with neural networks
with $\mathbf{u}$ representing the field variables of the physical system. From the perspective of dynamical systems, the state $\mathbf{u}$ is said to evolve on the attractor $A_{\mathcal{F}}$ of the underlying PDE system. For the following analysis, we define an attractor $A$ to satisfy the following conditions:
  1. 1.
    The attractor $A$ is forward invariant such that for any state $\mathbf{u}_{t=t_{0}}\in A$, the temporal evolution of this state is also in $A$ such that $\mathbf{u}_{t>t_{0}}\in A$. Thus, trajectories on $A$ cannot ”leave” this attractor and will further evolve within $A$.
  2. 2.
    The attractor has a basin of attraction $\mathcal{B}(A)$ in the neighborhood of $A$ for which all states $\mathbf{u}\in\mathcal{B}(A)$ converge to $A$ as $t\rightarrow\infty$. By this property, $A$ attracts neighboring states.
  3. 3.
    There is no subset of $A$ that satisfies the previous conditions.
The attractor can be seen as a subspace of the phase space of $\mathbf{u}$ containing only those realizations of $\mathbf{u}$ that are physical under Equation (1). For an initial condition in the basin of the attractor $\mathcal{B}(A)$, the solution will converge to states on $A_{\mathcal{F}}$ and further evolve within this attractor.
The learning tasks presented in the following sections assume that we have access to a highly-resolved ground truth solver that produces solution trajectories that are physical under Equation (1). Its trajectories follow the dynamics of a PDE $\mathcal{F}$ and thus converge to $A_{\mathcal{F}}$ for a reasonably chosen initial condition in $\mathcal{B}(A_{\mathcal{F}})$ and an adequate burn-in time. The resulting dataset of spatiotemporally discrete snapshots $\bar{U}=[\mathbf{\bar{u}}^{0}\dots\mathbf{\bar{u}}^{N}]$ lies strictly on this attractor such that $\bar{\mathbf{u}}^{t}\in A_{\mathcal{F}}$ for all $t$. Additionally, the physical system has a governing characteristic timescale $t_{c}$, where $t_{c}$ is dependent on the underlying PDE and its characteristic hyperparameters. The dataset size $N$ is typically large in comparison to this timescale.
2.2 Training Neural Simulators
We study Neural Simulator architectures that use neural networks for evolving discretized forms of PDE s. Simulating a system’s fully resolved dynamics becomes prohibitively expensive in many applications. A classic example is turbulent flows, where modeling approaches like large-eddy simulations often replace direct numerical simulations. In this context, Neural Simulators can applied on a reduced representation of the dynamical state. A mapping between the ground truth and its reduced representation leads to
$\tilde{\mathbf{u}}^{t}=R(\bar{\mathbf{u}}^{t}),\quad\forall\bar{\mathbf{u}}^{t}\in\bar{U}.$ (2)
It is important to note that the reduction is not strictly necessary. If one is interested in the full dynamics, the reduction $R$ becomes the identity. In our studies, the reduction depends on the physical system and will be specified for each system at a later stage.
The task of a Neural Simulator is to provide a time-stepping operation on discrete data frames in accordance with the PDE in (1). We denote the trajectory on which the Neural Simulator operates as $\mathbf{u}^{t}$. Due to modeling errors, $\mathbf{u}^{t}$ may not exactly fit $\tilde{\mathbf{u}}^{t}$. The Neural Simulator operation $g_{\theta}(\mathbf{u})$ on a sample from the dataset can then be denoted as
$\tilde{\mathbf{u}}^{t+1}\approx\mathbf{u}^{t+1}=g_{\theta}(\tilde{\mathbf{u}}^{t}),\quad$ (3)
on the dataset $\tilde{\mathbf{u}}^{t}\in\tilde{U}$, where $\theta$ represents the weights of the neural network. We want to stress that, at this point, we have not limited the structure of $g$ in any way. However, when specifying prediction and correction setups, we will differentiate the construction of $g$, even though both fit this general operation signature.
Prediction setups fully rely on a neural network to calculate the next timestep. Denoting the network as $f_{\theta}$ leads to
$\mathbf{u}^{t+1}=g_{\theta,p}(\mathbf{u}^{t})=f_{\theta}(\mathbf{u}^{t}).$ (4)
In a prediction configuration, the network fully replaces a numerical solver with the goal of improving its accuracy and performance. They’re also known as neural solvers or surrogates. Various previous works fit this description. It encompasses full state PDE predictions with a whole range of architectures including graph networks [79, 66], coarse-grained predictions of the full dynamics like in weather prediction [76, 39] as well as reduced order models or latent evolutions that operate on encoded spaces [104, 24]. We will focus on networks based on convolutional or graph message-passing operations. Details of the architectures are introduced with the physical system.
Correction setups are also concerned with time-evolving a discretized PDE but additionally include a numerical solver $\mathcal{S}$ that approximates the solution by computing only parts of a timestep. The underlying assumption is that the solver cannot capture the full dynamics of the PDE in (1). From a machine learning perspective, the numerical solver introduces a powerful prior and inductive bias. A timestep $g_{\theta}$ thus consists of the solver $\mathcal{S}$ and a neural network model $f_{\theta}$. The network complements the numerical solver such that
$\mathbf{u}^{t+1}=g_{\theta,c}(\mathbf{u}^{t};f_{\theta};\mathcal{S}).$ (5)
These setups represent a combination of numerical and neural architectures, potentially combining the advantages of both domains. Possible applications include closure and turbulence modeling in RANS or LES, coarse-graining, as well as dynamical system control setups. The formulation in (5) is quite general and fits many previous works. For instance, the correction networks used in previous work [93, 37] used the following more specific formulation
$\begin{split}\mathbf{u}^{*}&=\mathcal{S}(\mathbf{u}^{t})\\ \mathbf{u}^{t+1}&=f_{\theta}(\mathbf{u}^{*}).\end{split}$ (6)
We will also use this setup for our corrections. Different approaches exist. For instance in the context of LES, where the subgrid-scale terms are usually embedded into the filtered equation. This leads to
$\begin{split}\mathbf{f}&=f_{\theta}(\mathbf{u}^{t})\approx\nabla\cdot\tau_{\mathrm{sgs}}\\ \mathbf{u}^{t+1}&=\mathcal{S}(\mathbf{u}^{t},\mathbf{f}).\end{split}$ (7)
This closure-inspired formulation is more popular in the computational fluids community [83, 50]. In this context, training a neural network directly on ground truth values of $\tau_{\mathrm{sgs}}$ is referred to as a-priori training while training with respect to the solution state $\mathbf{u}$ is frequently called a-posteriori training [80].
For the prediction and correction setups from equation (6) and (4), respectively, the input and output vectors of the networks share the same solution space. While the learned mapping differs in prediction and correction cases, one does not necessarily need different architectures to train one or the other. Thus, we use different instances of identical neural network architectures in both setups. We use the same notation $f_{\theta}$ for both applications to stress this property. A representative computational chain for correction and prediction setups using this notation is visualized in Figure 2.